Today, you invest a lump sum amount in an equity fund that provides an 12% annual return. You would like to have $10,900 in 6 years to help with a down payment for a home. How much do you need to deposit today to reach your $10,900 goal? Do not round intermediate calculations. Round your answer to the nearest cent
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

10,900=P*(1.12)^6

P=10,900/(1.12)^6

which is equal to

=$5522.28(Approx).